# JANET: Joint Adaptive predictioN-region Estimation for Time-series

### v1

@@ 6.2.2 UK COVID-19 dataset
We evaluate JANET on a COVID-19 dataset of daily case counts from 380 UK regions (Stankeviciute et al., 2021; Sun and Yu, 2024). Table 2 presents the results. The task is to predict daily cases for the next 10 days based on the previous 100 days. Whilst the COVID case sequences from different regions are not independent, we anticipate at least approximate validity using our generalised ICP framework. JANET’s coverage is close to the desired significance level $\epsilon$ for most values of $K$. CF-RNN shows overcoverage with the basic RNN architecture but performs closer to the desired level with the EncDec architecture.
Table 2: Comparison of frequency of coverages (%) on the test set for $\epsilon=0.1$. The coverage values closer to 90% are highlighted in bold for every grouping of architecture (RNN, EncDec) on each dataset. Our method, JANET, is bolded. Both variants consistently achieve values close to 90% whereas CopulaCPTS often undercovers, MC-dropout significantly undercovers (failing completely for the COVID dataset), and CF-($\cdot$) often overcovers.
Method Particle1 Particle5 COVID-UK
$K=1$ MC-Dropout 79.34 43.40 0.00
CF-RNN 95.20 95.60 92.50
CopulaCPTS-RNN 89.60 90.40 85.00
JANET-RNN 85.80 89.80 88.75
CF-EncDec 98.80 92.40 87.50
CopulaCPTS-EncDec 85.60 86.40 78.75
JANET-EncDec 90.60 87.60 87.50
$K=2$ JANET-RNN 87.20 89.20 91.25
JANET-EncDec 90.20 88.00 87.50
$K=3$ JANET-RNN 87.00 89.00 91.25
JANET-EncDec 90.60 87.80 90.00
@@ 7 Discussion and future work
In this paper, we have formally extended the Generalised Conformal Prediction framework proposed by Chernozhukov et al. (2018) to the inductive conformalisation setting (Vovk et al., 2005). Building upon this foundation, we have introduced JANET (Joint Adaptive predictioN-region Estimation for Time-series), a comprehensive framework for constructing prediction regions in time series data. JANET is capable of producing prediction regions with marginal intervals that adapt to both the time horizon and the relevant historical context of the data. Notably, JANET effectively controls the $K$-FWER, a valuable feature particularly when dealing with long prediction horizons.
JANET includes several desirable properties: computational efficiency, as it requires only a single model training process; applicability to scenarios involving multiple independent time series, with exact validity guarantees; and approximate validity guarantees when a single time series is available. Furthermore, JANET is flexible enough to provide asymmetric or one-sided prediction regions, a capability not readily available in many existing methods. Whilst not directly implemented in this work, JANET also holds the potential for improved data efficiency by learning from permutations within individual time series when multiple independent time series are accessible, albeit at the cost of slightly relaxed coverage guarantees.
Looking towards future work, we envision several promising directions for extending JANET. One avenue involves exploring the extension of JANET to a cross-conformal setting (Vovk et al., 2005), which could offer gains in predictive efficiency at the expense of potentially weaker coverage guarantees. Additionally, we acknowledge a current limitation of JANET, which is its inability to create multiple disjoint prediction regions (Izbicki et al., 2021). Such disjoint regions could be denser than a single joint region, thereby providing more informative uncertainty estimates, particularly in cases with multimodal predictive distributions. We intend to address this limitation in future research.
Acknowledgments and Disclosure of Funding
@@ A.2.1 Assumptions for Theorem 1
Condition (1) states an ergodicity condition, that permuting the oracle conformity scores provides a meaningful approximation to the unconditional distribution of the oracle conformity score. This holds when a time series is strongly mixing ($\alpha$-mixing) Bradley (2007); Rio (2017); Bradley (2007); Rio (2017) using the permutation scheme discussed 4.2. Notably ARMA (Autoregressive and Moving Average) series with IID innovations are known to be $\alpha$-mixing (Bradley, 2007; Rio, 2017). Condition (2) bounds the discrepancy between the non-conformity scores and their oracle counterparts.
@@
Proof  The proof largely follows the proof of Generalised Conformal Prediction in Chernozhukov et al. (2018), adapted for Inductive Conformal Prediction. Since the second condition (bounds on the coverage probability) is implied by the first condition, it suffices to prove the first claim. Define the empirical distribution function of the non-conformity scores under randomization as:
@@
Theorem 2 (General Exact Validity)
@@
The proof follows from arguments for randomisation inference that can be found in Chernozhukov et al. (2018); Romano (1990).
Appendix B Training Details

### v2

@@ 5.2.2 UK COVID-19 dataset
We evaluate JANET on the UK COVID-19 dataset with daily case counts from 380 UK regions [35, 38]. Table 2 presents the results. The task is to predict daily cases for the next 10 days based on the previous 100 days. Whilst the COVID case sequences from different regions are not independent, we anticipate at least approximate validity using our generalised ICP framework. JANET’s coverage is close to the desired significance level $\epsilon$. CF-RNN shows overcoverage with the basic RNN architecture but performs closer to the desired level with the EncDec architecture.
Table 2: Comparison of coverage (%) and interval widths/areas on the test set for $\epsilon=0.1$. Coverage values closest to 90% are highlighted in bold for every grouping of architecture (RNN, EncDec) on each dataset. Narrower intervals are preferred. Coverages close to the desired significance level for $K=1$ are bolded. Results with * have been taken from existing works
Method Particle1 Particle5 UK COVID-19
Coverage Width Coverage Width Coverage Width
$K=1$ MC-Dropout 79.40 0.08 43.40 0.13 0.00 156
CF-RNN 95.20 0.12 95.60 0.25 92.50 921
CopulaCPTS-RNN 89.60 0.16 90.40 0.57 85.00 655
CAFHT-RNN 93.30 0.10 86.60 0.19 92.50 742
JANET-RNN(ours) 85.80 0.09 89.80 0.17 88.75 682
CF-EncDec 98.80 0.20 92.40 0.26 87.50 691
CopulaCPTS-EncDec 85.60 0.30 86.40 0.52 78.75 520
CAFHT-EncDec 94.60 0.22 93.00 0.22 92.50 610
JANET-EncDec(ours) 90.60 0.20 87.60 0.20 87.50 538
CCN-JPR* 91.00 - 89.00 - 87.00 -
$K=2$ JANET-RNN(ours) 87.20 0.06 89.20 0.12 91.25 593
JANET-EncDec(ours) 90.20 0.06 88.00 0.15 87.50 425
$K=3$ JANET-RNN(ours) 87.00 0.04 89.00 0.10 91.25 468
JANET-EncDec(ours) 90.60 0.05 87.80 0.13 90.00 361
6 Related work
Relaxing exchangeability
Recent research has shown a growing interest in adapting conformal prediction (CP) to non-exchangeable data. Early work by [41] explored relaxing the exchangeability assumption using Mondrian CP, which divides observations into exchangeable groups. [15] built upon this idea to share strengths between groups in hierarchical modelling. [39] and [30] addressed covariate and label shifts, respectively, by reweighting data based on likelihood ratios. Similarly, [28] and [9] applied reweighting for predictive inference in causal inference and survival analysis, while [18] focused on controlling covariate shift by statisticians. However, these methods address heterogeneity rather than the serial dependence found in time series and can not be applied to time series data.
One-step or multivariate prediction
[21] tackled distribution shifts in an online manner by adapting coverage levels based on comparisons of current coverage with desired coverage. [45] extended this work with online expert aggregation. [22] later introduced an expert selection scheme to guide update step sizes. These works typically require a gradient-based approach to learn a model that adapts to the coverage. [7] generalised pure conformal inference for dependent data, using fixed weights for recent training examples to account for distributional drift. [43] employed predictor ensembling, assuming exchangeability but with asymptotic guarantees. [10] leveraged randomisation inference [33] to generalise full conformal inference to serial dependence, achieving valid coverage under exchangeability and approximate validity with serial dependence. Notably, these methods primarily focus on single-step predictions for univariate series and can not be adapted . A similar extension to the inductive setting is provided in [13] for the functional time series setting. They randomly selected the time indices to form a calibration series, however, such an approach does not preserve the statistical properties of the time series. They also postulate different values of the predictor variable, which is not scalable to large time horizons. Additionally, the work is primarily based on applying to the time series and lacks formalisation as the generalised inductive conformal predictors. In contrast, we split the sequence into two such that we preserve the statistical properties of the time series; our method formally extends [10] to inductive conformal prediction and multi-step scenarios, including multivariate time series.
Other notable works, [12, 1, 44, 3] constructed prediction bands for multivariate functional data, functional surfaces, ellipsoidal regions for multivariate time series, prediction after skipping next time steps, respectively, but only for single-step predictions without an extension for multi-step predictions.
Multi-step prediction
[2] applied the jackknife method to RNN-based networks for multi-step prediction, with theoretical coverage of $1-2\epsilon$ at significance level $\epsilon$, instead of $1-\epsilon$ that we obtain and is desirable. [35] assumed conditional IID prediction errors in multi-output models, using Bonferroni correction [23] to achieve desired coverage. However, their approach can be overly conservative with increasing prediction horizons and may not hold when model assumptions are violated, and they require multi-output prediction models, in contrast to our approach that is model-agnostic. [11] utilised linear complementary programming and an additional dataset to optimise the parameters of a multi-step prediction error model, whereas JANET does not require any additional optimisation step or an additional dataset. In recent work, [38], building on [29], employed copulas to adjust for temporal dependencies, enabling multi-step and autoregressive predictions for multivariate time series filling in the deficit from other works. However, their method requires two calibration sets and gradient-based optimisation, making it data-inefficient, unlike JANET. Furthermore, unlike our proposed framework JANET, their approach requires multiple independent time series (just like the preceding works) and cannot adapt prediction regions based on historical context.
A concurrent work [46] primarily focussed on single-step prediction adapted [21] to account for heterogeneous trajectories in multiple time series settings. However, they require multiple calibration sets and do not show how to control $K-$familywise error when dealing with multi-step prediction. JANET, on the other hand, can control $K-$familywise error and does not require mutiple calibraiton steps. Another concurrent work [17] form multiple disjoint prediction regions, but their approach requires using generative models with access to density estimation, whereas JANET is model agnostic.
Non-conformal prediction regions
Beyond conformal methods, bootstrapping provides an alternative method for constructing joint prediction regions especially when one only has a single time series. [36] generates $B$ bootstraps and finds a heuristic-based prediction region from the bootstrapped predictive sequences, however, the method only provides have asymptotic guarantees. [42] creates JPRs that are asymptotically valid and can control the $K$-familywise error. [16] showed that model refitting in [36] can be avoided under certain assumptions but at the loss of asymptotic guarantees. However, their double bootstrapping method gets computationally expensive die to exponential inference from the trained models. However, given that these method are based on bootstrapping, they suffer from large computational cost that is infeasible when working with neural networks. JANET is based on inductive conformal prediction and has negligible computational cost in comparison to the bootstrapped methods. Our work can be seen as conformalisation of [42, 16] without relying on bootstrapping and with adaptive prediction regions that is not present in the prior works. Additionally, as a conformal method, JANET can be readily applied to time series classification tasks which is not apparent for bootstrap-based methods.
@@ 7 Discussion and future work
In this paper, we have formally extended the Generalised Conformal Prediction framework proposed by [10] to the inductive conformalisation setting [41]. Building upon this foundation, we have introduced JANET (Joint Adaptive predictioN-region Estimation for Time-series), a comprehensive framework for constructing prediction regions in time series data. JANET is capable of producing prediction regions with marginal intervals that adapt to both the time horizon and the relevant historical context of the data. Notably, JANET effectively controls the $K$-FWER, a valuable feature particularly when dealing with long prediction horizons.
JANET includes several desirable properties: computational efficiency, as it requires only a single model training process; applicability to scenarios involving multiple independent time series, with exact validity guarantees; and approximate validity guarantees when a single time series is available. Furthermore, JANET is flexible enough to provide asymmetric or one-sided prediction regions, a capability not readily available in many existing methods.
Looking towards future work, we envision several promising directions for extending JANET. One avenue involves exploring the extension of JANET to a cross-conformal setting [41], which could offer gains in predictive efficiency at the expense of potentially weaker coverage guarantees. Additionally, we acknowledge a current limitation of JANET, which is its inability to create multiple disjoint prediction regions [24]. Such disjoint regions could be denser than a single joint region, thereby providing more informative uncertainty estimates, particularly in cases with multimodal predictive distributions. We intend to address this limitation in future research.
Acknowledgment
@@ A.2.1 Assumptions for Theorem 1
Condition (1) states an ergodicity condition, that permuting the oracle conformity scores provides a meaningful approximation to the unconditional distribution of the oracle conformity score. This holds when a time series is strongly mixing ($\alpha$-mixing) [8, 32, 8, 32] using the permutation scheme discussed 3.2. Notably ARMA (Autoregressive and Moving Average) series with IID innovations are known to be $\alpha$-mixing [8, 32]. Condition (2) bounds the discrepancy between the non-conformity scores and their oracle counterparts.
@@
Proof.
The proof largely follows the proof of Generalised Conformal Prediction in [10], adapted for Inductive Conformal Prediction. Since the second condition (bounds on the coverage probability) is implied by the first condition, it suffices to prove the first claim. Define the empirical distribution function of the non-conformity scores under randomization as:
@@
Theorem 2 (General Exact Validity).
@@
The proof follows from arguments for randomisation inference that can be found in [10, 33].
Appendix B Computational Cost Analysis
In this section, we analsyse the computational costs for constructing prediction regions using different methods. Bonferroni and Scheffé-JPR require only a single run of the regression algorithm $\mathcal{A}$ (to fit $\hat{f}$ on the training set $Z_{\text{tr}}$), while bootstrap methods require $B$ separate runs—one for each bootstrap sample. Full conformal prediction (FCP)-based methods, like [10], in contrast, must train $\mathcal{A}$ many more times—once for each test point $X^{*}$ and each candidate label $y\in\mathcal{Y}$, where $\mathcal{Y}$ is a discretised grid of size $c$. The total number of model fits is therefore $(n_{\text{train}}+1)\cdot n_{\text{test}}\cdot c$.
In the case of JANET, only a single model fit is required as it is rooted in the Inductive Conformal Prediciton framework (ICP). The calibration cost depends on whether the task involves multiple time series or a single time series. For multiple time series, JANET evaluates one conformity score per calibration series—like standard ICP. For a single time series, however, JANET uses a permutation-based calibration strategy, computing one conformity score for each of the $d$ permutations of a single calibration sequence. Evaluation involves computing a $K-$max over $H$-dimensional residuals for each calibration permutation or series. For simplicity, we assume the model requires only one evaluation for $H$ forecasts into the future.
The Table 3 below summarises the model training and evaluation cost (ignoring constants), where all methods are trained on a dataset of size $n_{\text{tr}}$, and calibrated on $n_{\text{cal}}$ sequences (or permutations), for $n_{\text{test}}$ test examples. Table 3 compares computational cost (ignoring constants) for constructing prediction regions with different methods. The second column (Model training cost) counts the number of times the model $\hat{f}$ is trained. The third column (Model evaluation cost) accounts for computing conformity scores or predictions needed for calibration and test set evaluation. In most practical cases, the training cost dominates—for example, when $\hat{f}$ is a neural network, which is substantially more expensive to fit than to evaluate. Thus, methods like JANET, which require only a single model fit and relatively low calibration overhead, provide a practical alternative to bootstrap or full conformal methods, particularly in high-horizon or high-dimensional settings.
Table 3: Computational cost comparison of prediction region methods.
Method Model training cost Model evaluation cost
Bonferroni $1$ $n_{\text{test}}$
Scheffé $1$ $n_{\text{test}}$
Bootstrap-JPR $B^{2}\cdot n_{\text{test}}$ $B^{2}\cdot n_{\text{test}}$
Full conformal $(n_{\text{train}}+1)\cdot n_{\text{test}}\cdot c$ $(n_{\text{train}}+1)\cdot n_{\text{test}}\cdot c$
JANET (ours) $1$ $d+n_{\text{test}}$
Appendix C Additional experiments
C.1 Multiple Time Series
To further demonstrate JANET’s adaptability across architectures, we conducted additional experiments using two expressive forecasting models: (i) a lightweight Transformer, and (ii) a RealNVP-based normalising flow. The results are summarised in Table 4.
Transformer model.
We implemented a minimal Transformer encoder with one self-attention layer (1 layer, 2 attention heads) and a feedforward decoder. The conformal baselines (CF-Transformer, CopulaCPTS-Transformer) produce significantly wider regions than necessary, while still overshooting the target coverage of 90%. In contrast, JANET achieves near-nominal coverage with far more compact prediction regions, outperforming other methods in efficiency whilst also accounting uncertainty due to poor modelling.
Normalising flow model.
For probabilistic forecasting, we used a RealNVP-based model with three affine coupling layers. While these models produce full predictive distributions, their native prediction intervals are drastically miscalibrated. As shown in Table 4, the raw Flow model severely undercovers—sometimes catastrophically (e.g., 0% coverage on UK COVID-19). To test JANET’s robustness, we applied our method on the predicted means from the flow-model to compute the prediciton region.
The results confirm that JANET is able to recover valid joint prediction regions even from poorly calibrated probabilistic forecasts. Across all datasets and values of $K$, JANET-Flow achieves coverage close to or above 90% while maintaining narrow and interpretable prediction regions. This highlights JANET’s value as a post-hoc calibration framework that can correct for miscalibrated uncertainty estimates—whether they originate from deterministic architectures or full probabilistic models.
Table 4: Comparison of coverage (%) and interval widths/areas on the test set for $\epsilon=0.1$ using Transformer and Normalising Flow models. Coverage values closest to 90% are highlighted in bold. Narrower intervals are preferred.
Method Particle1 Particle5 UK COVID-19
Coverage Width Coverage Width Coverage Width
$K=1$ CF-Transformer 99.60 2.64 97.60 2.41 96.25 923
CopulaCPTS-Transformer 93.20 4.43 92.80 4.18 83.75 657
JANET-Transformer(ours) 88.40 1.65 89.40 2.04 87.50 744
Flow 19.20 0.27 15.80 0.29 0.00 1.8
JANET-Flow(ours) 91.40 0.44 91.80 0.80 87.5 20
$K=2$ JANET-Transformer(ours) 90.00 1.51 89.80 1.63 88.75 607
JANET-Flow(ours) 91.00 0.41 91.20 0.60 93.00 19
$K=3$ JANET-Transformer(ours) 90.20 1.48 89.80 1.50 88.75 537
JANET-Flow(ours) 91.00 0.40 91.00 0.44 91.00 18
Results.
Across both architectures, JANET achieves coverage rates close to the target $0.91-\epsilon=$ 0.9, while producing tighter regions than other baselines. Notably, JANET consistently improves the calibration of the probabilistic flow-based models, which often exhibit over- or under-coverage natively. These findings reinforce JANET’s robustness and generality across a wide range of forecasters, both deterministic and probabilistic.
C.2 Single Time Series Evaluation: Melbourne Minimum Temperature
To empirically evaluate JANET’s permutation-based calibration strategy in the single time series regime, we use the Melbourne Minimum Temperature dataset. [^1] Sourced from MachineLearningMastery.com (Jason Brownlee), available at https://github.com/jbrownlee/Datasets This dataset contains daily minimum temperatures in Melbourne, Australia, recorded from 1981 to 1990, for a total of 3652 observations. Two missing entries were imputed by averaging adjacent values to ensure compatibility with STL-based preprocessing.
We applied STL decomposition to extract and remove trend and seasonality components. The STL function identified yearly and weekly seasonalities. JANET’s permutation approach was applied to the residual (detrended and deseasonalised) series with the other components added later to the permuted residuals to obtain the set of calibration series.
Table 5: Empirical Out-of-Sample Coverages on the Melbourne Minimum Temperature Dataset ($\epsilon=0.2$, target coverage 80%), empirical width, and training times (minutes), numbers in the parenthesis refer to values of $K$. Bootstrap-JPR (Boot) and JANET perform similarly whereas Bonferroni (Bonf.) shows over coverage as usual and Scheffé undercovers. Bonf. and Scheffé can only be performed for $K=1$. The bootstrap methods take approximately 10-12 times as long as our method in wall-clock time in our implementations.
Cov. (%) 89 64 77 79 78 78 77 78
Width 11.83 6.55 8.56 8.33 6.21 6.02 6.03 5.99
Time (min) 10 10 8 91 119 119 11 119 11
Results.
Table 5 reports out-of-sample coverage, prediction width, and training time for JANET and several baselines. JANET achieves coverage close to the target level of $0.81-\epsilon=$ 0.8 across all values of $K$, with prediction regions that are consistently narrower than those produced by bootstrap-based joint prediction regions (Boot-JPR). Moreover, JANET is roughly 10–12$\times$ faster in wall-clock time due to its single-model training requirement.
Bonferroni-based intervals overcover, as expected, while Scheffé’s method undercovers significantly in this setting. In contrast, JANET offers a strong balance between coverage accuracy, interpretability, and computational efficiency—even when applied to a single, non-exchangeable time series where classical exchangeability assumptions may not hold.
C.2.1 Effect of Preprocessing on Coverage
JANET’s validity relies on approximate exchangeability of calibration residuals—an assumption that can be violated in the presence of strong non-stationarity due to trends or seasonal patterns. To understand how such violations affect coverage, we conducted an ablation study on the Melbourne Minimum Temperature dataset by systematically varying the level of preprocessing. Specifically, we compared JANET’s performance under three settings: (i) using the raw, non-stationary series, (ii) removing only the seasonal component using STL decomposition, and (iii) applying full STL decomposition to remove both trend and seasonality.
Table 6 summarises the coverage and width for $K=1,2,3$ at a fixed significance level of $\epsilon=0.2$. As expected, JANET undercovers when applied directly to the raw data, achieving only 71–73% coverage with relatively wide intervals. Removing seasonality improves both coverage and sharpness, while full STL decomposition results in near-nominal coverage (78–79%) and the narrowest regions.
Table 6: Coverage (%) and interval widths under different preprocessing strategies on the Melbourne Minimum Temperature dataset, evaluated at $K=1,2,3$ for $\epsilon=0.2$. We compare three cases: (i) using the raw non-stationary series, (ii) removing seasonality only, and (iii) full STL decomposition to remove both trend and seasonality. Coverage improves and width decreases as more structure is accounted for in the time series.
For $\epsilon=0.2$ K = 1 K = 2 K = 3
Raw series (no detrending or deseasoning)
Overall Coverage (%) 71 70 73
Overall Width 12.52 9.07 7.00
Deseasoned only (STL seasonal component removed)
Overall Coverage (%) 76 74 78
Overall Width 11.21 8.61 6.21
Full STL decomposition (trend + seasonality removed)
Overall Coverage (%) 79 78 78
Overall Width 8.33 6.02 5.99
These results confirm that violations of exchangeability due to non-stationarity can lead to predictable degradations in conformal validity. In the raw series setting, the poor coverage is expected—since the permutations used during calibration no longer correspond to a meaningful approximation of the original time series, thereby invalidating the underlying assumptions of JANET. However, we also observe that prediction intervals become substantially wider in this case, due to bigger residuals on the permuted series’, thus preventing a catastrophic collapse of coverage. In effect, although approximate validity is lost, JANET still produces non-trivial, cautious prediction regions. These findings reinforce the value of applying standard preprocessing pipelines (e.g., STL decomposition or differencing) to restore approximate exchangeability in real-world applications, and they motivate future work on the theoretical guarantees of conformal prediction under non-stationarity.
C.3 Time-Aware Weighting of Residuals
In some real-world applications, not all time steps in a forecast horizon are equally important. For example, earlier predictions in a sequence may be more actionable or reliable than those further ahead. In this section, we explore whether JANET can accommodate such asymmetries through time-aware calibration.
Although the default $K$-FWER construction treats all future time steps symmetrically, JANET’s non-conformity score can be flexibly adjusted to apply time-weighted scaling. Specifically, we reweight residuals from different segments of the forecast horizon when computing conformity scores, thereby biasing the prediction region construction toward earlier or later time steps.
We illustrate this idea using the RNN-based model on the Particle5 dataset. Table 7 compares three weighting schemes: (1) uniform weighting (default), (2) upweighting early time steps by downscaling their residuals by a factor of 1.5, and (3) upweighting late time steps by downscaling residuals in the last 12 steps by a factor of 1.5.
Table 7: Coverage (%) and interval widths under different residual weighting strategies on the Particle5 dataset, evaluated at $K=1,2,3$ for $\epsilon=0.1$. Errors are reported separately for the first and last 12 time steps of the prediction horizon. The final two rows in each block show how the fraction of errors shifts between early and late segments as a function of the weighting applied to residuals during calibration.
For $\epsilon=0.1$ K = 1 K = 2 K = 3
No scaling down (uniform)
Overall Coverage (%) 89.8 89.2 89
Overall Width 0.17 0.14 0.12
Fraction of Errors – First 12 steps 0.48 0.45 0.46
Fraction of Errors – Last 12 steps 0.52 0.55 0.54
Scale down residuals at early steps (×1.5 on first 12 steps)
Overall Coverage (%) 91.2 90.0 88.6
Overall Width 0.22 0.18 0.14
Fraction of Errors – First 12 steps 0.02 0.06 0.1
Fraction of Errors – Last 12 steps 0.98 0.94 0.9
Scale down residuals at late steps (×1.5 on last 12 steps)
Overall Coverage (%) 88.8 90.8 90.4
Overall Width 0.37 0.26 0.21
Fraction of Errors – First 12 steps 0.61 0.63 0.68
Fraction of Errors – Last 12 steps 0.39 0.37 0.32
As reported in Table 7, under uniform weighting, errors are distributed roughly evenly across the horizon. When earlier time steps are upweighted, JANET prioritises their coverage, significantly reducing the fraction of early-step errors—at the cost of more errors later in the horizon. Conversely, when later time steps are upweighted, the model shifts errors toward the beginning of the sequence. In all cases, JANET maintains approximate $K$-FWER coverage, demonstrating its flexibility in allocating uncertainty in a time-sensitive manner.
Appendix D Training Details